OpenStudy (haleholmes):

How do you simplify the following equation to make it easy to graph using a translation?

OpenStudy (haleholmes):



OpenStudy (dtan5457):

Wanna solve for X, right?

OpenStudy (haleholmes):

I don't know. It says make it easy to graph with translation.

OpenStudy (dtan5457):

too vague of a wording. do you just want to graph that?

OpenStudy (haleholmes):

Sure

OpenStudy (anonymous):

I would factor the 27 from both terms in 27x-81. This would give us \[\sqrt[3]{(27)(x-3)}\] The cube root of 27 is 3. So we would have this: \[3*\sqrt[3]{(x-3)}\] Now bring the -5 into the function. So this would be the simplified form I would use. \[3*\sqrt[3]{x-3}-5\] So the base function is going to be this: \[f(x) = \sqrt[3]{x}\] So looking at the simplified form, start with the base function and move the graph to the right 3 units in the positive x direction. Then expand the graph in the negative y and positive y directions by a multiple of three. Then move the graph down 5 units in the negative y direction.